Could you please help me set this word problem up and solve it? I am perplexed.
Paul flies his airplane 280 miles to the Moss County International Airport and picks up a load of bananas. He then flies back home. The trip to the airport was with the wind and the trip home was against the wind. The speed of the wind was 30 mph. If Paul's total flying time was 7 hours, what is the speed of Paul's plane in still air?

Answer by richwmiller(9135)

r*t=d
x+y=7
(r+30)*x=280
(r-30)*y=280
x=7/3 =2 1/3 hours with the wind
y=14/3= 4 2/3 hours against the wind
r=90 mph in still air.